OpenStudy (anonymous):

solve the separable differential equation: dy/dx= 2xycos(x^2) Someone please tell me how!! I know I have to integrate both sides but that's about all I know. solve the separable differential equation: dy/dx= 2xycos(x^2) Someone please tell me how!! I know I have to integrate both sides but that's about all I know. @Calculus1

OpenStudy (jamesj):

dy/dx= 2xycos(x^2) => dy/y = 2x cos(x^2) dx Now integrate.

OpenStudy (anonymous):

wait did you really bring the dx and y's to the other sides? I can now integrate the right side to be sin(x^2) but does integrating dy/y still cancel out effectively?

OpenStudy (anonymous):

orrr is it ln(y)?

OpenStudy (jamesj):

Yes, \[ \int \frac{dy}{y} = \ln y + C \] Make sure you add a constant to one side.

OpenStudy (anonymous):

I know this is just algebra stuff now, but that's what I always get confused on. How do you get from ln(y) + C = sin(x^2) to what I see as the answer: \[Ce^\sin(x^2)\] Is this correct: \[e ^{\ln(y) + C} = e ^{\sin(x^2)}\] and then e and ln(y) cross out to just being y, but how do I justify putting the C on the other side?

OpenStudy (jamesj):

exp[ln y + C] = y.e^c So y = A exp( sin x^2), where A = e^-C, that's how.

OpenStudy (anonymous):

Wait I didn't follow that at all. Could you re word possibly?

OpenStudy (jamesj):

\[e^{\ln y + C} = e^{\ln y}.e^C = y.e^C \] Yes?

OpenStudy (jamesj):

Hence if ln y + C = sin x^2, then y . e^C = e^(sin x^2) i.e., \[ y = e^{-C} e^{\sin x^2} \] \[ \ \ \ = A e^{\sin x^2} \] where A = e^-C

OpenStudy (anonymous):

ohhh wow you just replace e^-c with another variable for a constant. OK! well thank you soooooo much for your help! I'm teaching myself calculus this year because of a lack of course offerings at my school...so thank you for being my supplementary teacher!